It has been a long week here in the high bay for Kyle and myself -- but more so for Kyle. He has been on the late shift, getting in at about 3PM and working until 4, 5, or 6AM taking scans, babysitting fridge cycles, and analyzing data. It has taken its toll, and we're both looking forward to being done with this phase of the experiment.
Today was, of course, more scans, one of which seemed to be a waste of time (low signal-to-noise), the other quite reasonable. We tried to start taking another scan, but in the interests of Kyle's health and sanity, we are fridge cycling 'early' (which means Kyle will be in bed by 3AM!) and we will take the last of this calibration data tomorrow. After that, we break down the instrument (move the Ebert-Fastie, disconnect cables, etc) in preparation for the cryostat going on the gondola Saturday morning.
Britt and Michele finished Bemco'ing the power system, which passed with no problems (as expected based on the specs of the components). Kevin put Cynthia to work fleshing out our focal plane visualization tool, which will be nice to have soon! I also started work on a filter box to clean up the power coming out of the bolo system DC/DCs.